A boat cruises 36 miles against a 3 mph current and 36miles back in the direction of the same current. The round trip takes 5 hours. What is the speed of the boat in still water?
Answer by mananth(16949)

A boat cruises 36 miles against a 3 mph current and 36miles back in the direction of the same current. The round trip takes 5 hours. What is the speed of the boat in still water?
..
let speed of boat in still water be xmph
upstream speed = x-3
downstream speed = x+3
..
time = distance /speed
36/(x-3)+36/(x+3)=5
36(x+3)+36(x-3)=5(x^2-9)
36x+108+36x-108=5x^2-45
5x^2-72x-45=0
5x^2-75x+3x-45=0
5x(x-15)+3(x-15)=0
(5x+3)(x-15)=0
x= 15 mph --- speed of boat in still water
...
m.ananth@hotmail.ca